OpenStudy (anonymous):

One day, a person went to horse racing area, Instead of counting the number of human and horses, he instead counted 74 heads and 196 legs. Yet he knew the number of humans and horses there. How did he do it, and how many humans and horses are there? Help please. :) Thanx.

OpenStudy (anonymous):

x + y = 74 2x + 4y = 196 (2x + 4y) - (2x+ 2y) = 196 - 148 2y = 48 y = 24 x + (24) = 74 x = 74 - 24 x = 50 x = Humans y = Horses

OpenStudy (cwrw238):

let number of humans = x and number of horses now humans and horses each have only one head ( as if you didnt know!!) so x + y = 74 and also 2x + 4y = 196 solve these 2 equations simultaneously and you have your answer

OpenStudy (anonymous):

Thank you!!!